# SNR and Resource Adaptive Deep JSCC for Distributed IoT Image Classification

### v1

@@ Abstract
Index Terms:
Convolutional neural networks (CNN), joint source channel coding (JSCC), genetic algorithm, Random forest, image classification, Internet of Things (IoT).
@@ II System Model
The proposed network comprises a total of $B$ CNN blocks and is optimally split after $m$ blocks. Let the input image be denoted by $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, where each pixel value is normalized to the range $[-1,1]$ to prevent vanishing gradients. The normalized image $\mathbf{X}_{\text{norm}}$ is processed through the first $m$ CNN blocks located at the edge device. Each convolutional block performs the following operation:
@@
where $\mathbf{\hat{F}}_{m}$ is the recovered feature map, and $\mathbf{W}_{\text{recover}}$, $\mathbf{b}_{\text{recover}}$ are the the weight matrix and bias of the FC layer used for reconstruction. The remaining $B-m$ CNN blocks are executed at the server. The final feature map is projected into a 1D logit vector with dimension equal to the number of output classes. A global average pooling layer is applied, and a prediction is made for the input image. This concludes the forward pass.
@@ III-A Optimization Problem
where $k_{i}$ is the kernel size, $f_{\text{in}_{i}}$ and $f_{\text{out}_{i}}$ are the number of input and output filters, and $h_{i}$ and $w_{i}$ are the height and width of the feature map at layer $i$, respectively. For the fully connected layer mapping the feature space of $mth$ convolutional layer to the latent space, the number of FLOPs is given by:
@@ III-C Intelligent Genetic Algorithm
The goal is to determine the optimal values of $\Theta=\{f,k,l_{s},m\}$ that maximize image classification accuracy while satisfying constraints (9) and (10). We initialize a population of $V$ random configurations from the search space (Line $1$). For configurations already present in the offline dataset, FLOPs and accuracy values are directly retrieved (Lines $4-5$). Otherwise, they are predicted using the trained regressors $R_{F}$ and $R_{A}$ from Equations (14) and (15) (Lines $6-7$). Configurations violating the FLOPs constraint in (9) are discarded by assigning them a fitness score of $\infty$- ∞ (Lines $8-9$). Valid configurations are evaluated using a fitness function based on accuracy and resource utilization (Lines $10-11$). Each configuration $c$ receives a fitness tuple: $\text{fitness}(c)=(\left(A(c)+U(c),F_{\text{max}}-F(c),\right)$ where $U(c)=\max\left(0,\ \frac{F(c)-0.9\times F_{\text{max}}}{0.1\times F_{\text{max}}}\right)$.
@@
GA parameters: population size $V$, generations $G$, tournament size $T$, crossover probability $cxpb$, mutation probability $mutpb$
@@ 20      Perform tournament selection of size $T$;
22      Apply crossover with probability $cxpb$;
@@
24      Apply mutation with probability $mutpb$;
@@ 26      Update population $P$ with new configurations;
We apply a tournament selection strategy to promote strong yet diverse candidates (Line $12$). Specifically, $V$ overlapping subsets, each of size $T$, are formed from the current population. The top-performing configuration in each subset, based on the fitness score, is selected as a parent configuration. These parent configurations represent strong candidates from the current generation and are used to produce the next generation. Each pair of parents undergoes uniform crossover with probability $cxpb$, exchanging parameter values to promote exploration of the search space (Line $13$). Offspring configurations are then subjected to mutation with probability $mutpb$, where one parameter is randomly altered within its valid range (Line $14$). This process yields the next generation of configurations, which are evaluated and re-entered into the optimization cycle (Line $15$).
@@
From the final generation, we select the configuration with the highest predicted accuracy (Line $16$). In case of a tie, the configuration with the smallest FLOPs gap is chosen (Lines $17-18$). The resulting $\Theta_{\text{optimal}}$ is then used to configure the network in Fig. 1, ensuring optimal image classification performance within the given computational and SNR constraints.
@@ IV Performance Evaluation
Figure 2 shows the variation of the computational demand with the number of layers executed at the IoT device, using a fixed SNR of $-10\,\text{dB}$ and the corresponding optimal parameters. FLOPs are plotted on a logarithmic scale for better visual representation. As shown, the fixed-layer architecture from [9] incurs a constant FLOPs cost, making it unsuitable for scenarios with limited computational resources. At a low SNR of $-10\,\text{dB}$, the classification accuracy ranges from 42.33% to 66.15%, depending on the number of layers executed at the IoT device and the available computational budget.
@@
Figure 4 presents the accuracy of the proposed framework LAIGA compared to ADJSCC-l, evaluated across varying computational budgets and fixed SNR values. LAIGA consistently outperforms ADJSCC-l across all budget levels, primarily due to its computationally adaptive design. Even under low computational budgets, LAIGA enables full feature space propagation through the combined CNN layers on both the IoT device and the server. In contrast, ADJSCC-l relies on layer-wise transmissions only when refinement is required. The performance gain is especially notable at low SNR ($-10\,\text{dB}$) and tight computational budgets (1M–10M FLOPs), where our method achieves up to a 10% accuracy improvement. This gain is attributed to the SNR-robust training of the supervising dataset used during model optimization.

### v2

@@ II System Model
The proposed network comprises a total of $B$ CNN blocks and is optimally split after $m$ blocks. Let the input image be denoted by $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, where each pixel value is normalized to the range [1,1][-1,1] to prevent vanishing gradients. The normalized image $\mathbf{X}_{\text{norm}}$ is processed through the first $m$ CNN blocks located at the edge device. Each convolutional block performs the following operation:
@@
where $\mathbf{\hat{F}}_{m}$ is the recovered feature map, and $\mathbf{W}_{\text{recover}}$, $\mathbf{b}_{\text{recover}}$ are the the weight matrix and bias of the FC layer used for reconstruction. The remaining BmB-m CNN blocks are executed at the server. The final feature map is projected into a 1D logit vector with dimension equal to the number of output classes. A global average pooling layer is applied, and a prediction is made for the input image. This concludes the forward pass.
@@ III-A Optimization Problem
where $k_{i}$ is the kernel size, $f_{\text{in}_{i}}$ and $f_{\text{out}_{i}}$ are the number of input and output filters, and $h_{i}$ and $w_{i}$ are the height and width of the feature map at layer $i$, respectively. For the fully connected layer mapping the feature space of mthmth convolutional layer to the latent space, the number of FLOPs is given by:
@@ III-C Intelligent Genetic Algorithm
The goal is to determine the optimal values of $\Theta=\{f,k,l_{s},m\}$ that maximize image classification accuracy while satisfying constraints (9) and (10). We initialize a population of $V$ random configurations from the search space (Line $1$). For configurations already present in the offline dataset, FLOPs and accuracy values are directly retrieved (Lines 454-5). Otherwise, they are predicted using the trained regressors $R_{F}$ and $R_{A}$ from Equations (14) and (15) (Lines 676-7). Configurations violating the FLOPs constraint in (9) are discarded by assigning them a fitness score of -$\infty$ (Lines 898-9). Valid configurations are evaluated using a fitness function based on accuracy and resource utilization (Lines 101110-11). Each configuration $c$ receives a fitness tuple: $\text{fitness}(c)=(\left(A(c)+U(c),F_{\text{max}}-F(c),\right)$ where $U(c)=\max\left(0,\ \frac{F(c)-0.9\times F_{\text{max}}}{0.1\times F_{\text{max}}}\right)$.
@@
GA parameters: population size $V$, generations $G$, tournament size $T$, crossover probability cxpbcxpb, mutation probability mutpbmutpb
@@ 20  Perform tournament selection of size $T$;
22  Apply crossover with probability cxpbcxpb;
@@
24  Apply mutation with probability mutpbmutpb;
@@ 26  Update population $P$ with new configurations;
We apply a tournament selection strategy to promote strong yet diverse candidates (Line $12$). Specifically, $V$ overlapping subsets, each of size $T$, are formed from the current population. The top-performing configuration in each subset, based on the fitness score, is selected as a parent configuration. These parent configurations represent strong candidates from the current generation and are used to produce the next generation. Each pair of parents undergoes uniform crossover with probability cxpbcxpb, exchanging parameter values to promote exploration of the search space (Line $13$). Offspring configurations are then subjected to mutation with probability mutpbmutpb, where one parameter is randomly altered within its valid range (Line $14$). This process yields the next generation of configurations, which are evaluated and re-entered into the optimization cycle (Line $15$).
@@
From the final generation, we select the configuration with the highest predicted accuracy (Line $16$). In case of a tie, the configuration with the smallest FLOPs gap is chosen (Lines 171817-18). The resulting $\Theta_{\text{optimal}}$ is then used to configure the network in Fig. 1, ensuring optimal image classification performance within the given computational and SNR constraints.
@@ IV Performance Evaluation
Figure 2 shows the variation of the computational demand with the number of layers executed at the IoT device, using a fixed SNR of $10\,\text{dB}$ and the corresponding optimal parameters. FLOPs are plotted on a logarithmic scale for better visual representation. As shown, the fixed-layer architecture from [9] incurs a constant FLOPs cost, making it unsuitable for scenarios with limited computational resources. At a low SNR of $10\,\text{dB}$, the classification accuracy ranges from 42.33% to 66.15%, depending on the number of layers executed at the IoT device and the available computational budget.
@@
Figure 4 presents the accuracy of the proposed framework LAIGA compared to ADJSCC-l, evaluated across varying computational budgets and fixed SNR values. LAIGA consistently outperforms ADJSCC-l across all budget levels, primarily due to its computationally adaptive design. Even under low computational budgets, LAIGA enables full feature space propagation through the combined CNN layers on both the IoT device and the server. In contrast, ADJSCC-l relies on layer-wise transmissions only when refinement is required. The performance gain is especially notable at low SNR ($10\,\text{dB}$) and tight computational budgets (1M–10M FLOPs), where our method achieves up to a 10% accuracy improvement. This gain is attributed to the SNR-robust training of the supervising dataset used during model optimization.